OpenStudy (anonymous):

ok when i have a 90 degree angle = theta how would you graph it... would the hypotenuse connect the x axis and y axis or what?

OpenStudy (dumbcow):

you mean for a right triangle? yeah if you draw a line from y-axis to axis, that would form a right angle at the origin

OpenStudy (dumbcow):

*x-axis

OpenStudy (anonymous):

no it is a trig question so for quadrant angle but yes still a right angle

OpenStudy (dumbcow):

what exactly do you have to graph?

OpenStudy (anonymous):

\[\theta =3\pi \div2\]

OpenStudy (anonymous):

and i have to find sin cos tan... but i dont know if the graph should divide quadrant three like if it were a 45 degree angle or connectthe y axis to the x axis

OpenStudy (dumbcow):

ok use the unit circle 3pi/2 is an angle that goes from positive x-axis to negative y-axis draw line from positive x-axis to negative y-axis reference angle is 90 degrees sin 90 = 1 cos 90 = 0 sin refers to value of y-axis cos refers to value of x-axis since we are on negative y-axis sin 3pi/2 = -1 tan = sin/cos tan 3pi/2 = -1/0 = undefined

OpenStudy (anonymous):

ok